what is the formula of vibrational mode of energy ,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,,

Grade:11

1 Answers

Shresthahahaha
5 years ago
Vibrational energy,E=h(n+1/2)\nu=h(n+1/2)1/2\pi\sqrt{}(K/m)
here,n=quantum no,\nu=frequency
        m=mass
         h Planck’s constant